OpenStudy (anonymous):

The sum of four consecutive odd counting numbers is 80. If x represents the smallest number in the list, what are the four number? I tried x+(x+3)+(x+5)+(x+7)=80 and solved for x but it didn't come out to a whole number so I'm stuck. Help please?

OpenStudy (asnaseer):

if x is the smallest of these odd numbers, then the next consecutive odd number would be x+2 not x+3

OpenStudy (asnaseer):

take an example, next consecutive odd after 5 is 7 not 8

OpenStudy (anonymous):

@sunshnee8 Are you able to fix your mistake now?

OpenStudy (anonymous):

I still don't understand why it's x+2. It says consecutive odd counting numbers so I'm thinking 1,3,5,7,9...

OpenStudy (asnaseer):

yes that is a correct list of consecutive odd counting numbers - and notice each one is 2 away from the previous one.

OpenStudy (asnaseer):

1 3 = 1+2 5 = 3 + 2 ...

OpenStudy (anonymous):

I do notice that. I'm still trying to process the problem.

OpenStudy (asnaseer):

the mistake you made was to state that if x is the smallest of these numbers, then the next one would be x+3. it should be x+2 instead

OpenStudy (asnaseer):

then apply the same principal for the rest of them

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

A little. I found the numbers to be 17, 19,21,23. The problem is sometimes I don't know how I do it.

OpenStudy (asnaseer):

it just needs lots of practice - thats all. I'm sure you'll get the hang of these pretty soon. :)

OpenStudy (anonymous):

I've been struggling with word problems all my life. I hope I get to that point of understanding them soon haha. Thank you very much for your patience!

OpenStudy (asnaseer):

yw - I'm always glad to help - especially those who actually have a desire to learn - as you clearly do! :)

OpenStudy (anonymous):

Always wonderful to have people like you! Thanks!

OpenStudy (asnaseer):

thank you for your kind comment.